OpenStudy (anonymous):

A spherical snowball is melting at the rate of 2 cubic in. per minute and it's staying spherical! How fast is the radius of the snowball decreasing when the radius is 1/2 in?

OpenStudy (anonymous):

what is the area for a sphere? i forget

OpenStudy (anonymous):

\[\frac{3}{4}\pi r^3\] or something right?

OpenStudy (anonymous):

think so. ya

OpenStudy (anonymous):

well we can't guess, we need that one

OpenStudy (anonymous):

Will check. :)

OpenStudy (anonymous):

@satellite73 please fan me and ill fan u

OpenStudy (anonymous):

A = 4\pi r^2

OpenStudy (anonymous):

it is \[V=\frac{4}{3}\pi r^3\] take derivatives, you get \[V'=4\pi r^2 r'\] you know \(V'=\frac{1}{2}, r=\frac{1}{2}\) solve for\(r'\)

OpenStudy (anonymous):

ooops no sorry \(V'=2\)

OpenStudy (anonymous):

so \[2=4\times (\frac{1}{2})^2r'\] solve for \(r'\)

OpenStudy (anonymous):

it's one right?

OpenStudy (anonymous):

r = 2

jigglypuff314 (jigglypuff314):

oh right, r' isn't -2 because @satellite73 dropped a pi in that last comment

jigglypuff314 (jigglypuff314):

so -2 = 4 pi (1/2)^2 (r')

OpenStudy (anonymous):

one second jigles

OpenStudy (anonymous):

i don't know hot get the answer. im confused

jigglypuff314 (jigglypuff314):

well first try 4 (1/2)^2 1 4 * --- = 1 4 so you get -2 = pi (r') then divide both sides by pi

OpenStudy (anonymous):

-2pi?

jigglypuff314 (jigglypuff314):

divide pi

OpenStudy (anonymous):

by -2?

jigglypuff314 (jigglypuff314):

erm like -2/pi instead of -2pi

OpenStudy (anonymous):

−0.636619772368

jigglypuff314 (jigglypuff314):

yeah ... that xD

OpenStudy (anonymous):

but in fraction form?

jigglypuff314 (jigglypuff314):

mmm depends on what your teacher wants :)

OpenStudy (anonymous):

she wants fraction form i guess so its -2/pi for the final answer right?

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

ya, my darling is back. that's it. :)

OpenStudy (anonymous):

let r be the radius of snow ball at any time t \[then volume V=\frac{ 4 }{ 3 }\pi r ^{3}\] \[\frac{ dV }{dt }=\frac{ 4 }{ 3 } \pi *3 r ^{2}\frac{ dr }{dt }\] \[\frac{ dV }{dt }=-2 cubic inch\] when r=1/2 inch put the values \[-2=4 \pi 2 ^{2}\frac{ dr }{ dt }\] \[\frac{ dr }{dt }=\frac{ -2 }{16 \pi }=\frac{ -1 }{ 8 \pi } inch\] negative sign shows the radius is decreasing.

jigglypuff314 (jigglypuff314):

mmm you said r=1/2 but when you plugged it in it became "2^2" ? @surjithayer

OpenStudy (anonymous):

correction \[-2=4 \pi \left( \frac{ 1 }{ 2 } \right)^{2}\frac{ dr }{dt }\] \[\frac{ dr }{ dt }=\frac{ -1 }{\pi } inch\]

jigglypuff314 (jigglypuff314):

then how did -2 turn into -1?

OpenStudy (anonymous):

again wrong it should be -2/ pi inch

jigglypuff314 (jigglypuff314):

okay :) thanks for confirming :)

OpenStudy (anonymous):

thanks for correcting me.

OpenStudy (anonymous):

Jiggly Puff is the best!